I need help in understanding the steps to complete this problem
4x+6y=24
2x+3y=12

Answer by J2R2R(94)

To solve simultaneous equations you need two equations providing you with different information so you can use these differences to solve the problem.

The equations you have been given although different to some extent are telling you the same thing so aren’t really ‘different’ since one is a multiple of the other. So, you don’t have enough information to solve the equations.

4x+6y=24
2x+3y=12 contains the same information as 4x+6y=24 (above); 6x+9y=36; 8x+12y=48; etc.

This is easy to see when you have only two variables and two equations but when it gets more complicated with three and more variables and equations, sometimes one of the equations is a combination of the others and is therefore useless as you already have this information. This is beyond your question but I am explaining it to show that information contained in other equations you already have isn’t worth it and doesn’t help you solve your problem.